A 1.5 N stone is thrown vertically upward with an

initial speed of 15 m/s . Calculate

a. P.E at the maximum height

b. K.E at the instant it will strike the ground

c. P.E after 1 s

d. K.E after 1 s



1
Expert's answer
2022-03-25T15:02:28-0400

(a)

"PE_{\max}=KE_i=\frac{mv_i^2}{2}=\frac{1.5*15^2}{2*9.8}=17.2\:\rm J"

(b)

"KE_f=KE_i=17.2\:\rm J"

(c)

"PE=mgh=Wh=W(v_it-gt^2/2)\\
1.5*(15*1-9.8*1^2/2)=15.2\:\rm J"

(d)

"KE=KE_i-PE=17.2-15.2=5\:\rm J"